Suppose the position of an object moving in a straight line is given by

s left parenthesis t right parenthesis equals 7 t cubed minus t squared plus 5s(t)=7t3−t2+5.

Find the instantaneous velocity when

tequals=5
Answer #1

s(t)=7t^3-t^2+5
v(t)=ds(t)/dt=7*3*t^2-2*t
At t=5:
v(t)=7*3*5^2-2*5
v(t)=515

Instantaneous velocity=515